"Alex scored 13 points in the basketball game. This was 1/5 of the total points the team scored. Write and solve an equation to determine the total points (t) the points scored."
Answer by checkley77(12844)

13=1/5*x
13=x/5
x=5*13
x=65 ans. for the total points.